OpenStudy (anonymous):

Solve the system algebraically. 4x - 5y = 3 x = 3 - y What is the value of x? 9/7 24/14 33/14

OpenStudy (amistre64):

if you multiply the bottom by 5 and add the results you can solve for x

OpenStudy (anonymous):

?

OpenStudy (amistre64):

unless you get your hands dirty, youll never be able to do these on your own ....

OpenStudy (amistre64):

which part of the directions did you not get? multiply, 5, add, or solve ??

OpenStudy (amistre64):

4x - 5y = 3 x = 3 - y ; times 5 4x - 5y = 3 5x = 15 - 5y ; add em up ---------------- 9x -5y = 18 -5y ; and solve for x